Hi, I need help with this linear algebra problem. Thank you!

Let A = 0.7 0.1 0.3 0.9 Find P and D such that A = PDP-1, where D is a diagonal matrix.
Answer #1

A= 10.7 10.3 0.9 The characteristic equation of Ais. A11) = 72-face All + dela 72-1.62 +0.60 412) = 0 = 18 -1.61 +0.6 = 0 = 11503 : 1181-18) -0.39 + 0.1b = 0 let b = 3 then a=1 V₂ 3 i ] then 0 p={.:] =14 } ames=91T 3 A = Pop).